# Growing spines ad infinitum

### v1

@@ 5. An application: definability of henselian valuations
Let $k$ be a $t$-henselian field of characteristic 00. We first argue that we may replace $k$ with any $\mathcal{L}_{\textrm{ring}}$-elementary extension $k^{*}$: Assume that $k\preccurlyeq k^{*}\preccurlyeq k^{*}((H))$ holds for some non-trivial ordered abelian group $H$. We obtain elementary embeddings $k\preccurlyeq k^{*}((H))$ and $k\subseteq k((H))\preccurlyeq k^{*}((H))$, where the latter embedding is elementary by the Ax-Kochen/Ershov Theorem in equicharacteristic 00 [8, Theorem 6.17]. These embeddings imply that $k\preccurlyeq k((H))$ also holds.
In particular, by the paragraph above, we may assume that $k$ is $\aleph_{1}$-saturated in $\mathcal{L}_{\textrm{ring}}$. By [18, Lemma 3.3], $k$ admits a non-trivial henselian valuation $v$. Passing to an $\aleph_{1}$-saturated $\mathcal{L}_{\textrm{val}}$-elementary extension $(k^{*},v^{*})$ of $(k,v)$, we may further assume that $(k,v)$ is $\aleph_{1}$-saturated as a valued field. We next argue that $k$ admits a non-trivial henselian valuation $w$ of residue characteristic 00. Note that $v$ has a non-trivial coarsening $w$ of residue characteristic 00 if and only if $vk\supsetneq\braket{v(p)}_{convex}$ holds. That this inclusion is proper is ensured by $\aleph_{1}$-saturation of $(k,v)$. Thus, replacing $v$ by $w$, we may assume $\mathrm{char}(kv)=0$.
@@
Fact 5.3 (Beth definability theorem, [11, Thm. 6.6.4]).
@@
where $H_{i}$ are all copies of $H$ and with the obvious embeddings. By Tarski’s chain lemma [22, Theorem 2.1.4], $k$ is an elementary substructure of $k^{\prime}:=k((\sum_{i\in\mathbb{Z}}H))$. We show that there is a henselian valued field with residue $k^{\prime}$ whose valuation is not definable in $\mathcal{L}_{ring}$. The same property will hold for $k$ instead of $k^{\prime}$ by the $\preceq$-version of the Ax-Kochen/Ershov theorem ([8, Theorem 6.17]). Now consider $\Gamma:=\bigoplus_{i\in\mathbb{Z}}H$ and the Hahn field $K=k^{\prime}((\Gamma))$, together with the power series valuation $v$ with value group $\Gamma$. Consider the field $K^{*}=k^{\prime}((\Gamma^{\prime}))((\Gamma))$ with $\Gamma^{\prime}$ is another copy of $\Gamma$. One can notice, by playing Ehrenfeucht–Fraïssé games, that $\Gamma\oplus\Gamma^{\prime}$ is an elementary extension of $\Gamma$. The field $K^{*}$ admits the distinct henselian valuations
@@
Then we have by applying Ax-Kochen/Ershov [8, Theorem 6.17] once again
@@
The implication (1) $\implies$ (4) is essentially [6, Proposition 5.5]. Assume $k$ is not $t$-henselian, in particular $k$ is not separably closed. Consider the $\mathcal{L}_{\textrm{val}}$-theory $T$ which stipulates that any model $(K,v)$ is henselian with residue field $Kv\equiv k$ (in $\mathcal{L}_{\textrm{ring}}$) and the $\mathcal{L}_{\textrm{val}}$-formula $\phi(x)$ which asserts that $x\in\mathcal{O}_{v}$. Take $(K_{1},v_{1}),(K_{2},v_{2})\models T$ that have the same $\mathcal{L}_{\textrm{ring}}$-reduct, i.e., in particular $K_{1}=K_{2}$. Then both $v_{1}$ and $v_{2}$ are both henselian valuations with non-separably closed residue field, thus they are comparable, cf. [5, Theorem 4.4.2]. Assume for a contradiction that we have $\mathcal{O}_{v_{1}}\subsetneq\mathcal{O}_{v_{2}}$, then $v_{1}$ induces a non-trivial henselian valuation on $Kv_{2}$. But then $k$ is $t$-henselian, in contradiction to our assumption. Thus, we have $v_{1}=v_{2}$, and so Fact 5.3 implies that $v$ is uniformly $\emptyset$-definable in all models of $T$. ∎
@@
The crucial ingredient in Theorem 5.5 is Proposition 5.1, which fails in positive characteristic in general (cf. Remark 5.2). However, if $k$ is a perfect field, amending the proof of Theorem 5.5 (using [16, Theorem 7.1] instead of [8, Theorem 6.17]) shows we have the following chain of implications:
@@
Note that Theorem 5.5 may fail for imperfect fields: If $k$ is imperfect and admits no Galois extensions of degree divisible by $p$, then any henselian valuation with residue field $k$ is definable by [13, Proposition 3.6]. Note that such examples exist abundantly, e.g., $\mathbb{F}_{p}(t)^{\mathrm{sep}}$ is t-henselian (and in fact even henselian) but admits no proper Galois extensions of any degree.
Theorem 5.5 partially answers a question raised by Krapp, Kuhlmann, and Link ([15, §5.4]). Specifically, they inquire about identifying an optimal class $\mathcal{C}$ of fields such that for any henselian valued field with residue field in $\mathcal{C}$, the valuation is definable in the language of rings. Moreover, they ask the analogous question about identifying the subclass $\mathcal{C}_{0}$ of $\mathcal{C}$ such that such that for any henselian valued field with residue field in $\mathcal{C}$, the valuation is $\emptyset$-definable. Theorem 5.5 shows that restricting to the case of characteristic 00, the class of fields that are not $t$-henselian coincides exactly with both the classes $\mathcal{C}$ and $\mathcal{C}_{0}$.
Appendix. Iksrat’s Lemma and another proof of augmentability

### v3

@@ 5. An application: definability of henselian valuations
Let $k$ be a $t$-henselian field of characteristic 00. We first argue that we may replace $k$ with any $\mathcal{L}_{\textrm{ring}}$-elementary extension $k^{*}$: Assume that $k\preccurlyeq k^{*}\preccurlyeq k^{*}((H))$ holds for some non-trivial ordered abelian group $H$. We obtain elementary embeddings $k\preccurlyeq k^{*}((H))$ and $k\subseteq k((H))\preccurlyeq k^{*}((H))$, where the latter embedding is elementary by the Ax-Kochen/Ershov Theorem in equicharacteristic 00 [9, Theorem 6.17]. These embeddings imply that $k\preccurlyeq k((H))$ also holds.
In particular, by the paragraph above, we may assume that $k$ is $\aleph_{1}$-saturated in $\mathcal{L}_{\textrm{ring}}$. By [22, Lemma 3.3], $k$ admits a non-trivial henselian valuation $v$. Passing to an $\aleph_{1}$-saturated $\mathcal{L}_{\textrm{val}}$-elementary extension $(k^{*},v^{*})$ of $(k,v)$, we may further assume that $(k,v)$ is $\aleph_{1}$-saturated as a valued field. We next argue that $k$ admits a non-trivial henselian valuation $w$ of residue characteristic 00. Note that $v$ has a non-trivial coarsening $w$ of residue characteristic 00 if and only if $vk\supsetneq\braket{v(p)}_{convex}$ holds. That this inclusion is proper is ensured by $\aleph_{1}$-saturation of $(k,v)$. Thus, replacing $v$ by $w$, we may assume $\mathrm{char}(kv)=0$.
@@
Fact 5.3 (Beth definability theorem, [12, Thm. 6.6.4]).
@@
where $H_{i}$ are all copies of $H$ and with the obvious embeddings. By Tarski’s chain lemma [28, Theorem 2.1.4], $k$ is an elementary substructure of $k^{\prime}:=k((\sum_{i\in\mathbb{Z}}H))$. We show that there is a henselian valued field with residue $k^{\prime}$ whose valuation is not definable in $\mathcal{L}_{ring}$. The same property will hold for $k$ instead of $k^{\prime}$ by the $\preceq$-version of the Ax-Kochen/Ershov theorem ([9, Theorem 6.17]). Now consider $\Gamma:=\bigoplus_{i\in\mathbb{Z}}H$ and the Hahn field $K=k^{\prime}((\Gamma))$, together with the power series valuation $v$ with value group $\Gamma$. Consider the field $K^{*}=k^{\prime}((\Gamma^{\prime}))((\Gamma))$ with $\Gamma^{\prime}$ is another copy of $\Gamma$. Note that $\Gamma\oplus\Gamma^{\prime}$ is an elementary extension of $\Gamma$ (this holds, e.g., by playing Ehrenfeucht–Fraïssé games). The field $K^{*}$ admits the distinct henselian valuations
@@
Then we have by applying Ax-Kochen/Ershov [9, Theorem 6.17] once again
@@
The implication (1) $\implies$ (4) is essentially [7, Proposition 5.5]. Assume $k$ is not $t$-henselian, in particular $k$ is not separably closed. Consider the $\mathcal{L}_{\textrm{val}}$-theory $T$ which stipulates that any model $(K,v)$ is henselian with residue field $Kv\equiv k$ (in $\mathcal{L}_{\textrm{ring}}$) and the $\mathcal{L}_{\textrm{val}}$-formula $\phi(x)$ which asserts that $x\in\mathcal{O}_{v}$. Take $(K_{1},v_{1}),(K_{2},v_{2})\models T$ that have the same $\mathcal{L}_{\textrm{ring}}$-reduct, i.e., in particular $K_{1}=K_{2}$. Then both $v_{1}$ and $v_{2}$ are both henselian valuations with non-separably closed residue field, thus they are comparable, cf. [6, Theorem 4.4.2]. Assume for a contradiction that we have $\mathcal{O}_{v_{1}}\subsetneq\mathcal{O}_{v_{2}}$, then $v_{1}$ induces a non-trivial henselian valuation on $Kv_{2}$. But then $k$ is $t$-henselian, in contradiction to our assumption. Thus, we have $v_{1}=v_{2}$, and so Fact 5.3 implies that $v$ is uniformly $\emptyset$-definable in all models of $T$. ∎
@@
The crucial ingredient in Theorem 5.5 is Proposition 5.1, which fails in positive characteristic in general (cf. Remark 5.2). However, if $k$ is a perfect field, amending the proof of Theorem 5.5 (using [17, Theorem 7.1] instead of [9, Theorem 6.17]) shows we have the following chain of implications:
@@
Note that Theorem 5.5 may fail for imperfect fields: If $k$ is imperfect and admits no Galois extensions of degree divisible by $p$, then any henselian valuation with residue field $k$ is definable by [14, Proposition 3.6]. Note that such examples exist abundantly, e.g., $\mathbb{F}_{p}(t)^{\mathrm{sep}}$ is t-henselian (and in fact even henselian) but admits no proper Galois extensions of any degree.
Theorem 5.5 partially answers a question raised by Krapp, Kuhlmann, and Link ([16, §5.4]). Specifically, they inquire about identifying an optimal class $\mathcal{C}$ of fields such that for any henselian valued field with residue field in $\mathcal{C}$, the valuation is definable in the language of rings. Moreover, they ask the analogous question about identifying the subclass $\mathcal{C}_{0}$ of $\mathcal{C}$ such that such that for any henselian valued field with residue field in $\mathcal{C}$, the valuation is $\emptyset$-definable. Theorem 5.5 shows that restricting to the case of characteristic 00, the class of fields that are not $t$-henselian coincides exactly with both the classes $\mathcal{C}$ and $\mathcal{C}_{0}$.
A Iksrat’s Lemma and another proof of augmentability